john has 54 1/4 yard of material. If he needs to cut the cloth into pieces that are 3 1/12 yards long, how many pieces can he cut? how much material will be left over?
Answer by user_dude2008(1862)

(54 1/4)/(3 1/12) = (217/4)/(37/12) = (217/4)(12/37) = 651/37 = 17 22/37
Can cut 17 pieces with 22/37 yd left over